# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v3

@@ 4.1.2 LLM2CLIP Fine-tuning
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5 27.1
+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0 29.68
+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0 5.22
+Llama3-8B-TC(LLM2Vec) 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 86.5 83.3 28.42
+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2 29.46
+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 84.7 82.1 29.42
+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 87.8 84.3 29.96
Table 6: Comparison of various text encoders.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
@@
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2 73.5 70.0
       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0
+ Llama3-8B 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3 55.2 46.2
       ++ Dense Caption 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0
+ Llama3-8B-CC 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
       ++ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
Table 7: Ablation analysis of performance variations between short simple captions and long & dense captions using different text encoders. We conducted experiments on the EVA02 ViT-L/14 model, replacing its original text encoder with Jina-BERT (+ Jina-BERT), the original Llama3 (+ Llama3-8B), and Llama3 fine-tuned with Caption Contrastive learning (+ Llama3-8B-CC). Fine-tuning was performed on datasets containing only short simple captions and datasets with a 50% probability mix of long dense captions. The objective of this experiment was to isolate the effects of LLMs and long dense captions, and to demonstrate that the design of LLM2CLIP enhances the efficiency of dense caption utilization. The results with a red background indicate that the LLM without fine-tuning has a significant negative effect on CLIP.
Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
@@
100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5 84.2 80.6
75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2 87.3 84.1
50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0 87.0 83.8
0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
Table 8: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
Adaptor Layers Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
@@
0 91.3 79.9 64.0 50.2 97.1 97.7 88.1 90.2 81.0 86.1 84.3 80.8
1 92.1 80.6 65.9 51.8 97.8 98.0 90.9 91.8 84.3 87.9 86.2 82.0
4 92.1 79.9 66.0 52.2 97.4 98.0 93.0 92.1 85.0 88.0 86.7 82.0
Table 9: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different adaptor layer numbers. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. All experiments were conducted using OpenAI’s CLIP model with a ViT architecture at 224 resolution, trained on 15M data points.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
@@
Bidirectional Attention + MNTP + Sup. SimCLR 83.3 73.5 55.3 45.6 98.1 98.0 81.5 87.9 78.8 82.6 79.4 77.5
Bidirectional Attention + Sup. SimCLR 84.6 74.5 55.6 46.2 98.0 97.6 79.3 88.5 80.0 82.6 79.5 77.9
Sup. SimCLR 82.0 73.5 54.6 44.8 97.6 97.6 82.2 87.9 76.9 81.5 78.7 77.1
Table 10: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different methods. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. The experiments were conducted using the Llama 3.2 1B and the EVA02 ViT-L/14 model on 3M data. It can be observed that MNTP has no significant effect, while bidirectional attention provides a slight performance improvement.
4.2 Main Results
Evaluation Setting.
Our experiments assessed LLM2CLIP’s ability to enhance vanilla CLIP performance across multiple dimensions. We compared our approach against state-of-the-art (SOTA) models, including EVA [10] and OpenAI CLIP. Specifically, we evaluated CLIP ViT-B/16, ViT-L/14, and ViT-L/14-336, along with EVA02 ViT-B/16, EVA02 ViT-L/14, and EVA02 ViT-L/14-336, yielding a total of six comparison models.
4.2.1 Short- and Long-Text Retrieval
Evaluation Dataset.
@@ 4.2.1 Short- and Long-Text Retrieval
For short-text retrieval, we used the MSCOCO [23] 5K test set and the Flickr [43] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [5], the Urban1K [44] dataset, and the DOCCI [29] dataset.
@@
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [30], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
As shown in Table 2, LLM2CLIP fine-tuning yielded significant performance improvements, even surpassing models pretrained on large datasets like LAION2B. Compared to LongCLIP and JinaCLIP, our method achieved transformative gains, showcasing its efficiency in leveraging LLM knowledge to enhance pretrained CLIP models.

### v4

@@ 2 Related Works
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
LLM2Vec-Llama-3-8B 91.1 81.4 61.5 51.9 94.5 96.4 82.1 88.6 77.7 82.6 81.4 80.2
NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
Qwen2.5-0.5B-CC 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
Llama-3.2-1B-CC 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Llama-3-8B-CC 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
DeepSeek-R1-Distill-Llama-8B-CC 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
Llama-3.1-8B-CC 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
Table 6: Ablation for using different text encoder. BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs. The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
Methods(MLLM’s Caption Ratio) Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
       +LLM2CLIP (100%) 89.2 77.2 55.6 47.6 97.6 98.0 88.5 92.5 82.6 86.9 82.7 80.4
       +LLM2CLIP (75%) 92.0 81.3 61.5 52.0 97.7 97.8 85.1 92.2 84.0 87.7 84.1 82.2
       +LLM2CLIP (50%) 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
       +LLM2CLIP (25%) 91.1 81.8 61.8 53.0 97.3 97.2 83.4 90.6 83.9 86.9 83.5 81.9
       +LLM2CLIP (0%) 91.1 81.5 60.9 53.0 96.0 94.4 72.0 85.6 82.6 85.6 80.5 80.0
Table 7: Ablation study on the proportion of dense captions from different MLLMs. Here, we explore the impact of both genuine image captions and fake dense captions generated by MLLMs on the performance of  LLM2CLIP.
Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@
89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
Linear(×$\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
Linear(×$\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
Linear(×$\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Linear(×$\times 4$) Linear(×$\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
Transformer(×$\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
Transformer(×$\times 1$) Transformer(×$\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
[…]
3 Methods
3.1 Stage 1: LLM Caption Contrastive Fine-tuning
Traditionally, LLM output layers serve as classification heads producing text tokens. Prior works, such as LLM2Vec [2] and NV-EMBED-v2 [17], have explored converting LLMs into universal text encoders by applying additional designs and fine-tuning. Inspired by these approaches, we adapt similar concepts to the novel scenario of cross-modal pretraining. Our key design considerations encompass model architecture, training methods, and training data, summarized as follows:
Model Architecture:
  • Sentence Token Representation: We map the output layer features of the LLM to represent whole sentences. We considered two strategies: using the [EOS] token, which can attend to the entire sentence during pretraining, or employing average pooling across all output tokens. Empirically, average pooling performed better and is our default setting.
  • Bidirectional Attention: Given that the generative capability is unnecessary for our encoder-focused application, we remove the attention mask from the LLM, enabling bidirectional textual relationship modeling for enhanced comprehension.
  • Fine-tuning Parameters: To efficiently activate stronger textual comprehension capabilities within the output features of LLMs, we apply parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA).
  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability. We experimented with two adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, and (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs. However, neither approach significantly boosted CLIP performance during embedding fine-tuning. Consequently, we opted not to use adaptors during the LLM caption contrastive fine-tuning phase for simplicity.
Training Methods:
  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens. Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention. While effective in pure text tasks (as shown by LLM2Vec), MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
  • Textual Contrastive Learning: Following SimCSE, we employ textual contrastive learning to enhance feature separability, pulling positive samples closer and pushing negative samples apart in feature space. Two variants are possible: unsupervised (using dropout augmentation) and supervised (leveraging semantic pairs from annotated data). We default to supervised contrastive learning due to superior performance. Specifically, we generate positive pairs using two distinct captions from the same image, framed by a system prompt: ”Given a caption, retrieve a similar relevant caption.”
Training Data: We conducted experiments with Dreamlip [41]’s caption dataset, naturally containing multiple captions per image, making it ideal for supervised textual contrastive learning. Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings [29] into our contrastive training process. 89213
3.2 Stage 2: LLM2CLIP Post Fine-tuning
We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining. By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective post-training step. CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model. Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and determining the optimal approach to enhance CLIP performance is critical. Here, we elaborate on model architecture, training methods, and training data.
Model Architecture:
  • Fine-tuning Parameters and Adaptor Design: Training CLIP typically requires large-scale datasets and significant batch sizes, resulting in considerable computational costs. Therefore, we aim to employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this post-training stage. By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge. Given the large size of the LLM, we compare two fine-tuning strategies: (1) directly fine-tuning the LLM using LoRA, and (2) freezing the LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain the sentence token features. Method (2) offers distinct advantages, including completely eliminating gradient updates for the LLM, thus significantly reducing GPU memory consumption. Moreover, since the adaptor is placed after the LLM, we can even precompute and store text features through offline inference, thereby reducing the inference overhead from multiple epochs to just a single pass. Consequently, this approach eliminates the need to load the LLM into GPU memory during training, greatly accelerating experiment iteration and training time, and drastically reducing GPU memory footprint. By default, we adopt method (2) with an adaptor consisting of four linear blocks. Detailed efficiency analysis will be presented in Section 4.4.
Training Methods: We tested several approaches:
  • (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
  • (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
  • (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
  • (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
Extensive experiments (see Table 10) led us to adopt method (a) for its simplicity and effectiveness.
Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions. Datasets such as DreamLIP and Recaption, which leverage Share-Captioner [3], InstructBLIP [5] and LLaVA-1.5 [23] to generate multiple dense captions per image, can be effectively utilized. We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions. Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
Stage 1 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@ 3.2 Stage 2: LLM2CLIP Post Fine-tuning
Lora, AvgPool, Bidirectional, Supervise Simcse 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Lora, AvgPool, Bidirectional, Un-supervise Simcse 77.2 66.0 48.8 38.2 80.2 86.2 47.7 53.5 41.9 44.7 59.2 57.7
Lora, AvgPool, Bidirectional, MNTP 82.0 71.7 51.2 43.1 92.0 91.9 67.0 70.3 58.4 58.2 70.1 67.0
Lora, AvgPool, Bidirectional, MNTP, Supervise Simcse 88.2 78.3 59.6 48.9 95.5 95.1 79.7 86.1 75.3 77.7 79.7 77.2
Lora, AvgPool, Casual, Supervise Simcse 89.4 78.7 59.9 49.0 95.4 95.7 79.9 84.9 75.3 79.1 80.0 77.5
Lora, EOS, Bidirectional, Supervise Simcse 89.6 78.6 59.4 48.5 95.3 95.4 79.8 83.8 76.0 80.0 80.0 77.3
Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 85.5 75.5 54.1 45.3 92.0 93.5 72.3 75.5 66.4 66.9 74.1 71.3
Table 9: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.
Methods Training Loss Testing Text Encoder Flickr COCO ShareGPT4V Urban1k DOCCI Average
@@
CLIP CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune CL(CLIP-T, CLIP-V) CLIP-T 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
   +LLM2CLIP a) CL(LLM, CLIP) LLM 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
b) CL(LLM, CLIP-V)+ CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
LLM 90.1 81.3 61.3 52.4 97.1 97.4 85.6 91.1 83.8 86.9 83.6 81.8
c) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(CLIP-T, LLM) CLIP-T 89.0 77.9 58.5 47.7 89.1 88.8 67.0 75.5 66.3 70.7 74.0 72.1
LLM 90.8 80.5 61.5 51.8 97.1 97.7 85.9 90.3 83.4 86.7 83.7 81.4
d) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(Cat(CLIP-T, LLM), CLIP-V) CLIP-T 89.6 77.9 59.1 48.1 89.4 87.5 69.1 72.5 67.0 70.6 74.8 71.3
LLM 90.7 81.4 61.3 52.3 97.5 97.9 85.5 92.5 83.9 87.5 83.8 82.3
Cat(CLIP-T,LLM) 91.3 82.4 62.9 53.3 97.8 97.4 86.6 92.9 84.9 88.2 84.7 82.8
Table 10: LLM2CLIP Training Method Analysis. We experimented with various possibilities for post-training between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T). In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation. Items a–d correspond to the training methods section in Section 3.2. In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
Model Strategy Training Hour Batchsize Flickr COCO ShareGPT4V Urban1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
Llama-3.1-8B LLM LoRA 17 768 90.0 80.9 60.9 50.6 97.5 96.9 88.9 90.8 82.0 84.3 83.9 80.7
LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
Llama-3.1-8B-CC LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
LLM Frozen + Linear Adaptor 5.5 4096 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 93.2 82.2 64.2 54.0 98.2 98.3 87.8 93.1 86.2 88.8 85.9 83.3
Table 11: Efficiency study of LLM2CLIP. We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size. We report training hours, batch size, and retrieval accuracy for each experiment. During offloading, we precompute text captions offline into embeddings using the LLM. This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
4 Experiments
4.1 Default Experimental Settings
Default Settings for Caption Contrastive Fine-tuning We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention. No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding. MNTP is not used by default.
We utilize 30M DreamLIP caption data to perform caption contrastive fine-tuning on the LLM. Randomly sampling any two captions of an image from DreamLIP as positive pairs, we conduct supervised SimCSE training.
Default Settings for LLM2CLIP Cross-modal Post-training After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA 3.1 8B model to perform cross-modal contrastive learning post-training with the vision encoder. By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM. The design of the Adaptor is the same as FuseMix [35], using a simple inverted bottleneck MLP architecture. By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M. For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset. For the 3M setting, we only use the CC 3M subset. We summerize our default experiments setting and implementation details in Appendix A.
4.2 System Comparison
In the system comparison, we used the default settings from  Section 4.1 throughout. However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
4.2.1 Zero-Shot English Text-Image Retrieval
@@
Evaluation Dataset. For short-text retrieval, we used the MSCOCO [22] 5K test set and the Flickr [37] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [3], the Urban1K [39] dataset, and the DOCCI [26] dataset.
@@
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [27], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation. To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval. In Table Table 2, we compared OpenAI’s CLIP, EVA02, and Siglip2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoVit/14. Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02. For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8. Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed. Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
Training Data Volume Analysis for LLM2CLIP. In Table Table 2, we also conducted experiments with 3M, 15M, and 60M training data volumes on CLIP and EVA02 using ViT-L/14. We found that, whether at 224 or 336 resolution, more data consistently leads to better retrieval performance, for both long and short text retrieval tasks. Furthermore, we observed that the 3M data experiment on EVA02 is already sufficient to yield a noticeable improvement in long text retrieval tasks; however, no clear gain was observed in short text retrieval. We speculate that with a small data volume, the model first needs to compensate for the loss incurred by disrupting the original cross-modal space—a fact evidenced by the significant difference between the 3M and 15M data experiments.